OpenStudy (anonymous):

While standing on the roof of a building, a child tosses a tennis ball with an initial speed of 10 m/s at an angle of 40 degrees below the horizontal. The ball lands on the ground 3.3 s later. How tall is the building?

OpenStudy (anonymous):

Since you're looking for the vertical height, only concern yourself with vertical information.  Vy = 10sin(40) = 6.43 m/s Y = Yo + Vo*t - 0.5*g*t^2 Y = 0 m Yo = ? m Vo = 6.43 m/s t = 3.3 s g = 9.8 m/s^2 0 = Yo + (6.43)(3.3) - (0.5)(9.8)(3.3)^2 -Yo = 21.219 - 53.361 = -32.142 Yo = height = 32.142 m